# CCaaLF: Concurrency Control as a Learnable Function

### v3

@@ 5.3. Pipeline-Wait Action Optimization via Learned Conflict Graph
The Bayesian optimization technique in Section 5.2 can optimize $\mathcal{{F}}_{T}$ and $\mathcal{{F}}_{D}$. However, optimizing $\mathcal{{F}}_{P}$ is more challenging, especially under complex workloads, such as TPC-C, which involves many potential dependency cycles. Through extensive empirical analysis, we observe that fixing $\mathcal{{F}}_{T}$ and $\mathcal{{F}}_{D}$ while optimizing $\mathcal{{F}}_{P}$ leads to poor locality of the black-box function Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$) around the optimal solutions, i.e., a small change to $\mathcal{{F}}_{P}$ can lead to significant performance reduction due to new dependency cycles. Pipeline-wait actions, which are parameterized by $(w_{1},\dots,w_{n},expose)$, are designed to avoid such cycles because they lead to cascading aborts or deadlocks (DBLP:conf/sigmod/WangMCYCL16/ic3, ). As a result, the surrogate model trained via Bayesian optimization fails to approximate Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$) accurately, because the optimization inherently assumes smoothness and locality in the approximated function. Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ) discusses similar issues with Policy Gradient optimization which motivates it to adopt a genetic search algorithm.
@@
2 Function GraphReduction($branchFactor,mutateRate,K$):
@@ 3       $\mathcal{{V}}_{\textup{false}}\leftarrow$ a default function that returns false for all inputs
4       ${\mathcal{{F}}}_{best}\leftarrow$ the current best agent function.
@@ 15                               $Pop\leftarrow Pop-\{(\mathcal{{V}}_{e},\mathcal{{V}}_{o},score)\}$
22            $Pop\leftarrow$ GetTopK($Pop$)
23             if $Pop$ did not change in $MaxStep$ steps  then
@@ 29 Function Mutate($\mathcal{{V}}_{e},\mathcal{{V}}_{o},prob$):
31             Mutate $expose$ to true with $prob$ probability
@@
34             Mutate $optimistic$ to true with $prob$ probability
@@ 38             ${\mathcal{{F}}}_{best}\leftarrow{\mathcal{{F}}}_{new}$
To address the challenge of global optimization of $\mathcal{{F}}_{P}$, CCaaLF learns an optimal conflict graph from which the pipeline-wait actions can be derived, instead of learning these actions directly. This reformulation simplifies the learning process and improves optimization performance. A conflict graph is a static graph generated from the workload. In this graph, each type of transaction operation (e.g., a line of SQL) is a node. If two operations have potential conflicts (e.g., read-write conflicts), an edge is added between the corresponding nodes. CCaaLF uses SC-graph, a conflict graph widely adopted in existing CC algorithms (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/MuCZLL14/Rococo, ). These algorithms build a conflict graph offline and compute pipeline-wait actions by analyzing this graph prior to workload execution. However, they rely on static conflict graphs, which do not account for dynamic factors such as access skewness. We contend that the conflict graph needs to be dynamic, with conflict edges forming based on certain probabilities. For example, in the TPC-C workload, if there are many warehouses and few concurrent users, the update WAREHOUSE table access in the payment transaction is unlikely to conflict with the new order transaction’s read WAREHOUSE table access, and therefore this edge can be deleted. However, when there are few warehouses and many concurrent users, these accesses are more likely to conflict, therefore adding the conflict edge is beneficial.
@@
2 Function GetWaitActions($op,T,\mathcal{{V}}_{e},\mathcal{{V}}_{o},LastExposed$):
@@ 22             $res\leftarrow 0$
Algorithm 4 Pipeline wait actions
CCaaLF employs a graph reduction algorithm inspired by genetic search to quickly learn the optimal conflict graph. The algorithm starts with a large static conflict graph $G_{full}$, where each node represents a transaction operation $op$ from a transaction $T$, and an edge connecting two operations accessing the same table and least one operation is a write. The modifications of the conflict graph $G_{full}$ are defined by two functions, $\mathcal{{V}}_{e}$ and $\mathcal{{V}}_{o}$:
@@
2 Function UpdateWaitActions($\mathcal{{V}}_{e},\mathcal{{V}}_{o},{\mathcal{{F}}}_{best}$):
@@ 32                               $isOptimistic\leftarrow$ false
Algorithm 5 Updating pipeline wait actions
As outlined in Algorithm 3, the graph search algorithm starts with no conflict graph modifications, $\mathcal{{V}}_{e}=\mathcal{{V}}_{o}=\mathcal{{V}}_{\text{false}}$, then searches for better conflict graphs through mutations (lines 33 and lines 33). The algorithm maintains a population of the top $K$ best conflict graphs found during the search. The conflict graphs are evaluated by converting them into pipeline-wait actions (Algorithm 4), generating new agent functions with these pipeline-wait actions (Algorithm 5), and evaluating the corresponding agent functions. Specifically, given a conflict graph, we identify all potential dependency cycles and compute the required pipeline-wait actions to resolve these cycles (Algorithm 4 lines 44). After this, we update the pipeline-wait actions in the current best agent function to our calculated pipeline-wait actions and evaluate its performance score as the score for the conflict graph (Algorithm 3 line 3). As an agent function input may map to multiple nodes in the conflict graph, we select the longest wait action to prevent potential dependency cycles (Algorithm 5 lines 55). We adopt the same optimizations as IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ) to enforce the pipeline wait actions of the next operation before exposing dirty writes (Algorithm 1 lines 11), and to defer the pipeline waits of unexposed write operations until they are exposed (Algorithm 4 lines 44).
During optimization, the algorithm iteratively improves the conflict graphs by mutating the current graph population to better capture actual conflicts. Each mutation simplifies the graph by merging two consecutive nodes into a single node, or by removing conflict edges (Figure 6 and Algorithm 3). The mutation aims to eliminate unnecessary conflicts that will be resolved during the pipeline-wait actions calculation. Since the mutation reduces the number of nodes or the number of edges, the size of the graph reduces quickly towards the optimal conflict graph. The graph reduction search algorithm stops when no further mutations yield performance improvements (Algorithm 3 line 3), or when all mutations of the current population have been exhausted (line 3).
The graph reduction search algorithm complements the Bayesian optimization for $\mathcal{{F}}_{P}$. The former is efficient at exploring potentially optimal agent functions derived from conflict graphs, which Bayesian optimization alone may struggle to explore. CCaaLF still relies on Bayesian optimization to approximate high-performance but sparsely distributed agent functions. For instance, in the YCSB workload, we observe that a no-pipeline-wait policy can outperform learned pipeline-wait actions due to reduced computational costs associated with conflict detection and dependency tracking. Such an agent function cannot be identified through graph reduction search, but it can be found quickly by Bayesian optimization within minutes.
@@
5.4. Agent Function Optimization Pipeline
@@
Given a time budget, it is important to prioritize the optimization of actions that are more likely to yield significant performance improvements. CCaaLF uses an optimization pipeline that transfers knowledge from one optimization stage to the next. It allocates the time budget efficiently across these stages. Recall that the agent function $\mathcal{{F}}$ is decomposed into three sub-functions $\mathcal{{F}}_{P}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{T}$ (see Section 5.1). The optimization pipeline consists of four stages, each targeting a subset of actions while keeping the remaining actions unchanged from the best-seen agent function. Each stage leverages the previously evaluated agent functions and their corresponding performance scores as initial training data for the stage’s optimizer.
@@
The first stage optimizes both $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ using the graph reduction algorithm. These two sub-functions are optimized first because they lead to the largest performance gains, e.g., with improvements of nearly 80% in the TPC-C workload. $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ can be optimized simultaneously, as actions no detection and cut-edge mutations both assume no conflicts around an operation. This stage starts with the same initial configuration as that of IC3, in which all conflict detection actions are set to detect critical. The graph reduction search iteratively refines the conflict graph. When an agent function input $s$ maps to conflict graph nodes that are entirely disconnected from others, it indicates that database operations in state $s$ rarely encounter conflicts. In such cases, we set the conflict detection action to no detect to reduce transaction execution costs (Algorithm 5 lines 55). The graph population size is set to $K=4$, enabling quick convergence to an efficient agent function. The stage ends when no further mutations yield performance improvements, typically within 10 minutes due to the monotonically reducing nature of the mutation process.
@@
The second stage optimizes $\mathcal{{F}}_{T}$ using Bayesian optimization. By this stage, most parameters in $\mathcal{{F}}_{T}$ are continuous and exhibit good locality. This stage ends when no further performance improvements are observed after several consecutive evaluations (e.g., 20 evaluations). The third stage is the same as the first stage, but with a larger population size ($K=8$). This stage allows for deeper exploration of the conflict graph, in order to refine the conflict graph and agent functions. Finally, the fourth stage fine-tunes all the parameters of $\mathcal{{F}}$ using Bayesian optimization. This stage continues until the time budget is exhausted.
5.5. Feaure Selector Optimization
@@
We formulate the feature selector as a function that maps a feature selection strategy to the performance score achieved by CCaaLF with the corresponding selected features. Each feature selection strategy is represented as a vector, with each element indicating whether a corresponding feature is included and specifying its transformation type (linear, square root, or logarithmic). The output of the feature selection strategy is a set of input features for the agent function $\mathcal{F}$, which is subsequently used during the agent function optimization. Similar to the agent function optimization described in Section 5.2, Bayesian optimization is employed to iteratively propose new feature selection strategies, evaluate them, and refine the surrogate model. This process identifies the set of features that maximizes the performance score.
@@
To balance effectiveness and efficiency, the feature selection strategy is optimized to maximize the performance improvement achieved through function optimization within a user-defined time budget. Note that selecting too few features can overly constrain the search space of agent functions, leading to suboptimal solutions, whereas selecting too many features can expand the search space excessively, making the optimization computationally expensive. Therefore, we balance the two factors by incorporating the time budget into the performance score. We allocate a default budget of one hour and run the feature selector optimization for one day to determine the feature selection strategies used in our experiments. Because the optimization is costly, we pre-optimize a selector that performs well under the default settings. The user can opt to run the optimization to further improve the default selector’s performance. The optimized feature selectors and their configurations are available in our code repository (CCaaLF-impl, ).
@@ 6. Performance Evaluation
In this section, we evaluate the performance of CCaaLF and compare it with state-of-the-art concurrency control (CC) algorithms under diverse workloads. We implement CCaaLF in C++ and integrate it into Silo (DBLP:conf/sosp/TuZKLM13/silo, ), a widely used database system for evaluating CC algorithms. Specifically, we add the following components to the database engine: a collector that captures transaction-related features, a feature selector that refines the captured features, an in-database cache that stores the learned agent functions, and a monitor that tracks system performance for function optimization and workload drift detection. We modify the concurrency control code in the database engine to integrate the logic presented in Algorithm 1. We further implement a script that processes optimization requests and training data from the database. Our code is available at (CCaaLF-impl, ).
@@ 6.1. Experimental Setup
The experiments are conducted on a 24-core Intel(R) Xeon(R) Gold 5317 server with 123GB of memory. The CPU runs at a base frequency of 800 MHz, and a maximum turbo frequency of 3.6 GHz.
Execution. CCaaLF executes transactions in one of the following two modes: stored procedure mode (DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/SuCDAX17/Tebaldi, ; DBLP:conf/icde/Su0Z21/C3, ; DBLP:conf/sigmod/GuoWYY21/bamboo, ; DBLP:conf/sigmod/LinC0OTW16/LEAP, ; DBLP:journals/vldb/YaoZLOX18/DGCC, ) and interactive transaction mode (DBLP:journals/tocs/ZhangSSKP18/TAPIR, ; DBLP:journals/tocs/CorbettDEFFFGGHHHKKLLMMNQRRSSTWW13/spanner, ; DBLP:conf/sigmod/PavloCZ12/skew_aware, ; DBLP:journals/pvldb/YuBPDS14/1000_core_cc, ).
@@
  • Stored procedure mode. This mode represents how stored procedure transactions are executed by the database engine. In particular, all the transaction’s operations are known before execution. Therefore, concurrency control algorithms can use features such as transaction type and access ID. The transaction only returns results to the user after the transaction commits or aborts, which means it can read uncommitted data during execution and employ database-controlled partial retries.
@@
  • Interactive transaction mode. This mode represents how interactive, user-driven transactions are executed. In particular, the transaction’s operations (or SQL queries) are issued and executed individually. The user waits for the response of each operation before issuing the next. Because the operations are not known, concurrency control algorithms cannot rely on features such as transaction type and access ID. Furthermore, since the results are returned to the user immediately after each operation, CC algorithms cannot perform actions such as partial retry and reading dirty data from uncommitted transactions.
@@
Workloads. We conduct experiments on the widely used YCSB, TPC-C, and TPC-E workloads. We change three types of workload settings for evaluation: access pattern, contention level, and system load. The workloads and their settings are described below.
@@
  • YCSB-extended. We extend YCSB to group multiple read/write operations into a single transaction, and simulate varying access patterns by changing hotspot positions. Specifically, we fix the list of read and write operations for all transactions, representing stable operations found in real-world applications  (DBLP:conf/osdi/WangDWCW0021/polyjuice, ; DBLP:conf/sigmod/WangMCYCL16/ic3, ). We then vary the positions of hotspots, where hotspot accesses follow a highly skewed distribution (Zipfian factor $\theta=1$), while other accesses follow a uniform distribution across all keys (Zipfian factor $\theta=0$). For this workload, we use 16 threads.
@@
  • TPC-C. We use the standard TPC-C workload, which consists of five transaction types: two read-only transactions and three read-write transactions. This workload has complex inter-transaction dependencies, making it harder to optimize than YCSB-extended. We vary the contention level by changing the number of warehouses (fewer warehouses mean higher contention). We vary the system load by changing the number of threads (more threads mean a higher system load). For this workload, the default setting is 16 threads and one warehouse. For a fair comparison, we follow Polyjuice to include all five types of TPC-C transactions. We learn the concurrency control for three types, i.e., payment, new order, and delivery, while excluding the remaining two types, as they are read-only transactions.
  • TPC-E. We employ a TPC-E variant used in (DBLP:conf/sigmod/WangMCYCL16/ic3, ; DBLP:conf/osdi/WangDWCW0021/polyjuice, ), which includes three read-write transactions trade order, trade update, and market feed. Compared to the TPC-C and YCSB-extended workloads, this workload has more operations per transaction, resulting in a larger search space for learning-based algorithms. We vary the contention level by changing the access skewness, which is controlled via the update operations to the SECURITY table (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), following a Zipfian distribution. The change of the system load setting in the TPC-E workload is the same as that of the TPC-C workload.
We note that both CCaaLF and other learned concurrency control algorithms adapt to workload drifts by first detecting the drifts and then reoptimizing their algorithms. Specifically, the agent functions (in CCaaLF) or policies (in Polyjuice) are retrained when workload drifts are detected.
@@
Baselines. We compare CCaaLF against five state-of-the-art concurrency control algorithms: Silo (DBLP:conf/sosp/TuZKLM13/silo, ), Polyjuice (DBLP:conf/osdi/WangDWCW0021/polyjuice, ), IC3 (DBLP:conf/sigmod/WangMCYCL16/ic3, ), 2PL (2PL_no_wait, ), and CormCC (DBLP:conf/usenix/TangE18/CormCC, ). For the first four baselines, we use the implementations in (DBLP:conf/osdi/WangDWCW0021/polyjuice, ). For CormCC, we follow the approach in Polyjuice and report the best performance between 2PL and Silo.

### v4

@@ 5.2. Pipeline-Wait Action Optimization via Learned Conflict Graph
The Bayesian optimization technique in Section 5.1 can effectively optimize $\mathcal{{F}}_{T}$ and $\mathcal{{F}}_{D}$, but optimizing $\mathcal{{F}}_{P}$ is more challenging, especially for complex workloads like TPC-C that involve many potential dependency cycles. We observe empirically that during the optimization of Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$), the function exhibits poor locality with respect to $\mathcal{{F}}_{P}$ around the optimal solutions. Even minor changes to $\mathcal{{F}}_{P}$ can significantly reduce performance by introducing new dependency cycles that lead to cascading aborts or deadlocks (Wang et al., 2016). As a result, the surrogate model trained via Bayesian optimization fails to approximate Score($\mathcal{{F}}_{T}$, $\mathcal{{F}}_{D}$, $\mathcal{{F}}_{P}$) accurately, because the optimization inherently assumes smoothness and locality in the approximated function.
@@
2Function GraphReduction($branchFactor,mutateRate,K$):
@@ 3  $\mathcal{{V}}_{\textup{false}}\leftarrow$ a default function that returns false for all inputs
4  ${\mathcal{{F}}}_{best}\leftarrow$ the current best function.
@@ 18              $Pop\leftarrow Pop\cup\{(\mathcal{{V}}_{e}^{\prime},\mathcal{{V}}_{o}^{\prime},score^{\prime})\}$
22    $Pop\leftarrow$ GetTopK(PopPop)
23     if PopPop did not change in MaxStepMaxStep steps then
@@ 29 Function Mutate($\mathcal{{V}}_{e},\mathcal{{V}}_{o},prob$):
31     Mutate exposeexpose to true with probprob probability
@@
34     Mutate optimisticoptimistic to true with probprob probability
@@ 38     ${\mathcal{{F}}}_{best}\leftarrow{\mathcal{{F}}}_{new}$
To address the challenge of global optimization of $\mathcal{{F}}_{P}$, NeurCC learns an optimal conflict graph instead of learning these actions directly. The optimal conflict graph is the learned conflict graph from which the pipeline-wait actions that maximize the performance score are derived. This reformulation simplifies the learning process and improves optimization performance. A conflict graph is a graph generated from the workload, where each type of transaction access (e.g., a line of SQL statement) is represented as a node. If two operations have potential conflicts (e.g., read-write conflicts), an edge is added between the corresponding nodes. NeurCC uses SC-graph, a conflict graph widely adopted in existing CC algorithms (Wang et al., 2021, 2016; Mu et al., 2014). These algorithms construct a static conflict graph and compute pipeline-wait actions by analyzing this graph before workload execution. However, they rely on static conflict graphs, which do not account for dynamic factors such as access skewness. We contend that the conflict graph should be dynamic, with conflict edges added based on certain probabilities. For example, in the TPC-C workload, when there are many warehouses and few concurrent users, the payment transaction’s update WAREHOUSE table access is unlikely to conflict with the new order transaction’s read WAREHOUSE table access, and therefore this edge can be deleted. However, when there are few warehouses and many concurrent users, these accesses are more likely to conflict, therefore adding the conflict edge is beneficial.
@@
2Function GetWaitActions($op,T,\mathcal{{V}}_{e},\mathcal{{V}}_{o},LastExposed$):
@@ 22     $res\leftarrow 0$
Algorithm 4 Fetching pipeline-wait actions
NeurCC employs a graph reduction search algorithm inspired by genetic search to quickly learn the optimal conflict graph. The algorithm starts with a large static conflict graph $G_{full}$, where each node represents a transaction operation $op$ from a transaction $T$, and an edge connecting two operations accessing the same table and least one operation is a write. The modifications of the conflict graph $G_{full}$ are defined by two functions, $\mathcal{{V}}_{e}$ and $\mathcal{{V}}_{o}$:
@@
2Function UpdateWaitActions($\mathcal{{V}}_{e},\mathcal{{V}}_{o},{\mathcal{{F}}}_{best}$):
@@ 32              $isOptimistic\leftarrow$ false
Algorithm 5 Updating pipeline-wait actions
@@
Two types of graph reduction mutations.
As outlined in Algorithm 3, the graph search algorithm starts with no conflict graph modifications, $\mathcal{{V}}_{e}=\mathcal{{V}}_{o}=\mathcal{{V}}_{\text{false}}$, then searches for better conflict graphs through mutations (lines 33 and lines 33). The algorithm maintains a population of the top $K$ best conflict graphs found during the search. The conflict graphs are evaluated by converting them into pipeline-wait actions (Algorithm 4), generating new functions with these pipeline-wait actions (Algorithm 5), and evaluating the corresponding functions. Specifically, given a conflict graph, we identify all potential dependency cycles and compute the required pipeline-wait actions to resolve these cycles (Algorithm 4 lines 44). After this, we update the pipeline-wait actions in the current best function to our calculated pipeline-wait actions and evaluate its performance score as the score for the conflict graph (Algorithm 3 line 3). As a function input may map to multiple nodes in the conflict graph, we select the longest wait action to prevent potential dependency cycles (Algorithm 5 lines 55). We adopt the same optimizations as IC3 (Wang et al., 2016) to enforce the pipeline-wait actions of the next operation before exposing dirty writes (Algorithm 1 lines 11), and to defer the pipeline waits of unexposed write operations until they are exposed (Algorithm 4 lines 44).
During optimization, the algorithm iteratively improves the conflict graphs by mutating the current graphs to better capture actual conflicts. Each mutation simplifies the graph by edge removal or by merging consecutive nodes (as illustrated in Figure 6 and Algorithm 3). The mutation aims to eliminate unnecessary conflicts that will be resolved during the pipeline-wait actions calculation. Since the mutation reduces the number of nodes or the number of edges, the size of the graph reduces quickly towards the optimal conflict graph. The graph reduction search algorithm stops when no further mutations yield performance improvements (Algorithm 3 line 3), or when all mutations of the current population have been exhausted (line 3).
The graph reduction search algorithm complements the Bayesian optimization for $\mathcal{{F}}_{P}$. The former efficiently explores functions derived from conflict graphs, while the latter is effective at approximating high-performance but sparsely distributed functions. For instance, in the YCSB workload, we observe that a no-pipeline-wait policy can outperform learned pipeline-wait actions due to reduced computational costs associated with conflict detection and dependency tracking. Such a function cannot be identified through graph reduction search, but it can be found quickly by Bayesian optimization within minutes.
5.3. Function Optimization Pipeline
@@
Given a limited time budget, it is important to prioritize the optimization of actions that are more likely to yield significant performance improvements. NeurCC uses an optimization pipeline that transfers knowledge from one optimization stage to the next. It allocates the time budget efficiently across these stages. Recall that the function $\mathcal{{F}}$ is decomposed into three sub-functions $\mathcal{{F}}_{P}$, $\mathcal{{F}}_{D}$, and $\mathcal{{F}}_{T}$ (see Section 5). The optimization pipeline consists of four stages, each targeting a subset of actions while keeping the remaining actions unchanged from the best-seen function. Each stage leverages the previously evaluated functions and their corresponding performance scores as initial training data for the stage’s optimizer.
@@
The first stage optimizes both $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ using the graph reduction search algorithm. These two sub-functions are optimized first because they lead to the largest performance gains, e.g., with improvements of nearly 80% in the TPC-C workload. $\mathcal{{F}}_{P}$ and $\mathcal{{F}}_{D}$ can be optimized simultaneously, as the no detection action and edge removal mutation both assume no conflicts around an operation. This stage starts with the same initial configuration as that of IC3, in which all conflict detection actions are set to detect critical. The graph reduction search iteratively refines the conflict graph. When a function input $\mathbf{{s}}$ maps to conflict graph nodes that are entirely disconnected from others, it indicates that database operations in state $\mathbf{{s}}$ rarely encounter conflicts. In such cases, we set the conflict detection action to no detect to reduce transaction execution costs (Algorithm 5 lines 55). We set $K=4$ to enable quick convergence to an efficient function. The stage ends when no further mutations yield performance improvements, typically within 10 minutes due to the monotonically reducing nature of the mutation process.
@@
The second stage optimizes $\mathcal{{F}}_{T}$ using Bayesian optimization. At this stage, most parameters in $\mathcal{{F}}_{T}$ are continuous and exhibit good locality. This stage ends when no further performance improvements are observed after several consecutive evaluations (e.g., 20 evaluations). The third stage is the same as the first stage but with $K=8$, allowing for deeper exploration of the conflict graph, in order to refine the conflict graph and functions. Finally, the fourth stage fine-tunes all the parameters of $\mathcal{{F}}$ using Bayesian optimization. This stage runs until the time budget is exhausted.
5.4. Feature Selector Optimization
@@
We define the feature selector $\mathcal{{E}}$ as a function mapping a feature selection strategy to the system performance achieved with the selected features. Each feature selection strategy is represented as a vector, with each element indicating whether a corresponding feature is included and specifying its transformation type (linear, square root, or logarithmic). The output of $\mathcal{{E}}$ is a set of input features for $\mathcal{F}$, which is subsequently used during the function optimization. As in Section 5.1, NeurCC uses Bayesian optimization to iteratively propose new feature selection strategies, evaluate them, and refine the surrogate model.
@@
To balance effectiveness and efficiency, the feature selection strategy is optimized to maximize the performance improvement achieved through function optimization within a user-defined time budget. Note that selecting too few features can overly constrain the search space of functions, leading to suboptimal solutions, whereas selecting too many features can expand the search space excessively, making the optimization computationally expensive. Therefore, we balance the two factors by incorporating the time budget into the performance. In our experiments, we run the feature selector optimization for one day, using a function optimization budget of one hour for each feature selection evaluation.
The optimization is costly and is performed only once when the system is deployed. Subsequent changes in system load, contention, or access patterns do not automatically trigger re-optimization. Instead, NeurCC relies on function optimization to handle such workload drifts. However, users can explicitly run the optimization to fine-tune the default selector for specific workloads. Our feature selectors and their configurations are available in the code repository (6).
6. Deployment
To support practical adoption, we provide a complete guideline for deploying NeurCC in a new environment. The process consists of the following four steps.
Transaction Annotation (Optional). When transactions are executed in the stored procedure mode, the user first assigns a transaction type to each transaction and a unique access ID to every SQL statement. The transaction type can be assigned per stored procedure, while the access ID can be derived from the statement order (e.g., the line number of the API invocation within the procedure (Wang et al., 2021; Nguyen et al., 2025)). Next, the user constructs a large static conflict graph as the initial graph, where each node represents a type of transaction access (an access ID), and edges connect accesses on the same table if at least one is a write.
Optimization Pipeline Setup. Next, the user configures the optimization pipeline for function optimization. This includes specifying the number of optimization steps, selecting the optimization algorithm for each step (e.g., graph reduction search or Bayesian optimization), defining termination conditions for Bayesian optimization steps, and setting the population size for graph reduction search. Our default optimization pipeline is described in Section 5.3. The user may also select an initial function according to their needs. By default, NeurCC initializes with IC3 in stored procedure mode and 2PL in interactive transaction mode.
Feature Selector Optimization. After configuring the optimization pipeline, the user runs the feature selector optimization once to reduce the state space and accelerate future function optimization. As discussed in Section 5.4, this step runs for one day by default. Because it is costly, it is performed only once during initial deployment. Users may also choose to manually trigger this optimization periodically (e.g., once per month) or on demand (e.g., when migrating to a new machine).
Execution and Function Optimization. Finally, the user runs NeurCC with the optimized feature selector and the configured optimization pipeline. During execution, the system continuously monitors throughput and automatically detects workload drift (Section 4.3). When workload drift is detected, it automatically triggers function optimization to produce a better function for subsequent transaction processing.
@@ 7. Performance Evaluation
In this section, we evaluate the performance of NeurCC and compare it with state-of-the-art concurrency control (CC) algorithms under diverse workloads. We implement NeurCC in C++ and integrate it into Silo (Tu et al., 2013), a widely used database system for evaluating CC algorithms. Specifically, we add the following components to the database engine: a collector that captures transaction-related features, a feature selector that selects influential features, an in-database cache that stores the learned functions, and an optimizer that tracks system performance for function optimization and workload drift detection. We integrate Algorithm 1 into the concurrency control code of the database engine. We implement a script that processes optimization requests from the database. The code is available at (6).
@@ 7.1. Experimental Setup
All experiments are conducted on a 24-core Intel(R) Xeon(R) Gold 5317 server with 128GB of memory. The CPU runs at a base frequency of 800 MHz, and a maximum turbo frequency of 3.6 GHz.
Execution modes. NeurCC executes transactions in one of the two modes: stored procedure (Wang et al., 2016, 2021; Su et al., 2017, 2021; Guo et al., 2021; Lin et al., 2016; Yao et al., 2018; Nguyen et al., 2025), and interactive transaction (Zhang et al., 2017; Corbett et al., 2013; Pavlo et al., 2012; Yu et al., 2014).
@@
  • Stored procedure mode. This mode represents how stored procedure transactions are executed by the database engine. In particular, all the transaction’s statements are known before execution. As a result, concurrency control algorithms can use transaction type and access ID as features. The former can be assigned to each stored procedure, and the latter can be derived from the statement order, e.g., the line number of the API invocation within the procedure (Wang et al., 2021). The transaction only returns results to the user after the transaction commits or aborts, which means it can read uncommitted data during execution and perform partial retries.
@@
  • Interactive transaction mode. This mode represents how interactive, user-driven transactions are executed. In particular, the transaction’s statements (SQL queries) are issued and executed individually. The user waits for the response of each statement before issuing the next. Because the operations are not known in advance, concurrency control algorithms cannot rely on features such as transaction type and access ID. Furthermore, since the results are returned to the user immediately after each operation, the algorithms cannot perform partial retry and dirty read.
@@
Workloads. We conduct experiments on the widely used YCSB and TPC-C workloads. We change three types of workload settings for evaluation: access pattern, contention level, and system load. The workloads and their settings are described below.
@@
  • YCSB-extended. We extend YCSB to group multiple read/write operations into a single transaction, and simulate varying access patterns by changing hotspot positions. Specifically, we fix the list of read and write operations for all transactions, representing stable operations found in real-world applications  (Wang et al., 2021, 2016). We then vary the positions of hotspots, where hotspot accesses follow a skewed distribution (Zipfian factor $\theta=1$), while other accesses follow a uniform distribution across all keys (Zipfian factor $\theta=0$). For this workload, we use 16 threads.
@@
  • TPC-C. We use the standard TPC-C workload, which consists of five transaction types: two read-only transactions and three read-write transactions. This workload has complex inter-transaction dependencies, making it harder to optimize than YCSB-extended. We vary the contention level by changing the number of warehouses (fewer warehouses mean higher contention). We vary the system load by changing the number of threads (more threads mean a higher system load). For this workload, the default setting is 16 threads and one warehouse. For a fair comparison, we follow Polyjuice to include all five types of TPC-C transactions. We learn the concurrency control actions for three types, i.e., payment, new order, and delivery, while excluding the remaining two types, as they are read-only transactions.
@@
Baselines. We compare NeurCC against five state-of-the-art concurrency control algorithms: Silo (Tu et al., 2013), Polyjuice (Wang et al., 2021), IC3 (Wang et al., 2016), 2PL (Chesnais et al., 1983), and CormCC (Tang and Elmore, 2018). For the first four baselines, we use the implementations in (Wang et al., 2021). For CormCC, we follow the approach in Polyjuice and report the best performance between 2PL and Silo. We refer to this baseline as $CormCC^{*}$.